OpenStudy (anonymous):

Circle A has center of (0, 4) and a radius of 6, and circle B has a center of (-3, 5) and a radius of 24. What steps will help show that circle A is similar to circle B?

OpenStudy (anonymous):

@Preetha @sleepyjess

OpenStudy (anonymous):

I need someone else's opinion I'm stuck

OpenStudy (anonymous):

Aren't all circles simialr?

OpenStudy (sleepyjess):

I never thought I would be in the same tag as preetha lol

OpenStudy (anonymous):

similar

OpenStudy (sleepyjess):

And I agree with @eliassaab

OpenStudy (sleepyjess):

@amistre64 Can you confirm this?

OpenStudy (amistre64):

the steps to show that one circle is similar to another is not simply stating that they are similar ...

OpenStudy (anonymous):

Yes, but I have to pick one of the multiple choice

OpenStudy (amistre64):

how do we show things are similar? what is your definition?

OpenStudy (amistre64):

one definition may be: 2 things are similar if one can be transformed into the other only by dilation

OpenStudy (anonymous):

Translate circle A using the rule (x+3, y−1). Rotate circle A 180° about the center. Dilate circle A by a scale factor of 4. Reflect circle A over the line y=x.

OpenStudy (amistre64):

well, that might be a technically bad definition on my part, since moving objects from one place to another is valid since the structure isnt changed.

OpenStudy (amistre64):

if we have 2 circles, and move them so that they center about the same point; then we can dilate one circle into the other in a general way to prove that all circles are similar to each other

OpenStudy (amistre64):

no need to rotate a circle ... but other than that, it seems fair to me

OpenStudy (amistre64):

if we can determine a center of dilation, then moving the circles becomes redundant as well

OpenStudy (amistre64):



OpenStudy (anonymous):

No those are four different choices lol 1. Translate circle A using the rule (x+3, y−1). 3. Dilate circle A by a scale factor of 4. 4. Reflect circle A over the line y=x.

OpenStudy (anonymous):

So you think it's dilate by four as well? It's not translate circle A?

OpenStudy (amistre64):

if we are given to only one process, dilation is the most important yes. reflecting does nothing for us, and translating doesnt prove similarity, but does help in getting them to a more accessible dilation point is all. its dilation that is the important process to me

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

if we ignore centers, then dilating a radius if 6 into a radius of 24 to get congruent cirlces is acceptable yes C1 = 2pi (r1) C2 = 2pi (r2) in order for a dilation of C1 to equal C2 d C1 = C2 d 2pi r1 = 2pi r2 d r1 = r2 d = r2/r1 therefore d = 24/6 in this case